Calculate the amount of energy in kilojoules absorbed or released when 17.8 grams of H2O is formed during the reaction shown below. Use the Heat of Formation Table to calculate ΔHrxn. Balance the equation before beginning the calculation.

HCl(aq) + Ca(OH)2(aq) → CaCl2(aq) + H2O(l)



Please help!
1
Expert's answer
2017-04-26T08:05:48-0400
Solution:
The balanced chemical equation is:
2HCl(aq) + Ca(OH)2(aq) → CaCl2(aq) + 2H2O(l)
According to Hess’s law, we can calculate the value of ΔHrxn, using formula:
ΔHrxn = (ΔH2980 (CaCl2) + 2·ΔH2980(H2O)) – (2·ΔH2980(HCl) + ΔH2980(Ca(OH)2)
Using the Heat of Formation Table:
ΔH2980 (CaCl2) = - 785.8 kJ/mol;
ΔH2980(H2O) = - 285.84 kJ/mol;
ΔH2980(HCl) = - 92.30 kJ/mol;
ΔH2980(Ca(OH)2 = -986.2 kJ/mol.
Thus:
ΔHrxn = (- 785.8 + 2·(- 285.84)) – (2·(- 92.30) + (-986.2)) = -186.68 (kJ/mol)
If ΔHrxn ˂ 0, the chemical reaction is exothermic and energy released.
The number of moles of water equals:
n (H2O) = m(H2O)/M(H2O) = 17.8/18 = 0.99 (g)
Make a proportion:
186.68 kJ released for 1 mole
X kJ released for 0.99 moles
Where:
X = (0.99∙186.68)/1 = 184.81 (kJ) = Q

Answer: Q = 184.81 kJ.